OpenStudy (anonymous):

During a recent trip to the store, a car's velocity went from 0 to 60 mph in 20 seconds. What is the average acceleration of the car in mile per hour per hour

OpenStudy (amistre64):

id assume it to be the slope of the line created by the points (0,0) and (20,60)

OpenStudy (amistre64):

lol ... they appear to have a censoring database up and working

OpenStudy (anonymous):

Ages since I did these, is that per hour or per minute?

OpenStudy (anonymous):

i converted the given time to hours. so it is per hour

OpenStudy (anonymous):

At least I get the 0 to 60 part,lol

OpenStudy (anonymous):

U have 20 seconds/60

OpenStudy (anonymous):

= 1/3 hour.

OpenStudy (anonymous):

yes correct...

OpenStudy (anonymous):

Ah mph....

OpenStudy (anonymous):

Yes, I remember now, have to check dimensions...

OpenStudy (anonymous):

NOOOOOOO !! i made a boo-boo it has to be divided by 3600 to convert to hour !!!

OpenStudy (anonymous):

Ah, I wasn't imagining things then..

OpenStudy (anonymous):

10,800, right'

OpenStudy (anonymous):

Average acceleration is the rate at which velocity changes. Average acceleration is the change in velocity divided by an elapsed time. change in velocity = 60 mph - 0mph = 60 mph time taken for this change = 20 seconds = 20/3600 h = 1/180 h 60 180 average acceleration = ------ mph^-2 = 60 * ---- = 10800 mph^-2 1/180 1 That's WOW acceleration.......

OpenStudy (anonymous):

Did the driver survive this acceleration??????? How did the car achieve this acceleration???

OpenStudy (anonymous):

Yes, I guess as a rate it goes up a lot over an hour.

OpenStudy (anonymous):

I think not feasible for an hour...

OpenStudy (anonymous):

....☺

OpenStudy (anonymous):

Oh well, we gave him an answer, anyway...

OpenStudy (anonymous):

yeah, true......

OpenStudy (anonymous):

u deserve a medal for pointing out the error in my calculation....

OpenStudy (anonymous):

Thanks I have to do some reading up on these things, I used to quite like mechanics/physics things, but years ago...

OpenStudy (anonymous):

yeah, it is an interesting area to look into...........